OpenStudy (anonymous):

Find a plane containing the point (5,4,-6) and the line of intersection of the planes -x+2y+8z=-57 and 5x-4y+7z=-27

OpenStudy (loser66):

The intersection of 2 planes ( if there is) is a line, what is it? ( I mean the equation of that line?)

OpenStudy (anonymous):

The question does not given any other information

OpenStudy (loser66):

sure, but we have a way to find it out, :) you are supposed to know it.

OpenStudy (anonymous):

Would I need to find the normal vector? So the cross product of those two equations?

OpenStudy (loser66):

@sourwing please help

OpenStudy (loser66):

ok, he left. I don't have time, but let me try.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Since I tried solving the cross product and got 46i+47j-6k

OpenStudy (loser66):

let say P1 , P2 are plane 1, 2, so normal vector of P1 is n1 = < -1,2,8> and n2 = <5,-4,7> ok?

OpenStudy (anonymous):

Okay

OpenStudy (loser66):

then n1 x n2 =?

OpenStudy (anonymous):

I got 46i+47j-6k

OpenStudy (anonymous):

Then do I need to make x=0 and solve for y and z?

OpenStudy (loser66):

That is the intersection line, and you need 2 points in that line, right? but if you want to go on your way, you let x =0 for the equations of 2 planes, not this line I mean -x+2y+8z=-57 and 5x -4y +7z = 27 let x =0 for both , solve for y and z

OpenStudy (loser66):

what do you have?

OpenStudy (anonymous):

Hold on...

OpenStudy (loser66):

2y + 8z = -57 -4y +7z = 27 so, y = -615/46 z= -87/23

OpenStudy (loser66):

I am sorry, I don't have time, cannot wait for you :)

OpenStudy (anonymous):

No worries!

OpenStudy (loser66):

so, you have a point (0, -615/46, -87/23)

OpenStudy (anonymous):

And that would be my (x0,y0,z0) point?

OpenStudy (anonymous):

And (46,47,-6) would be my (a,b,c) points?

OpenStudy (loser66):

therefore, the equation of the intersection line is x = 46t y =( -615/46) + 47t z = (-87/23) -6t now, just let t =0 to get point A (...,...,..) and then t =1 to get point B (...,...,...) accompanied with the given point (5,4,-6) 3points , so you can write down the equation of the required plane, right?

OpenStudy (anonymous):

Got it. I'll try to go through that process then.

OpenStudy (loser66):

ok, good luck, sorry for not letting you work on your problem, I need my time to write my lab report. try to help because I started but sourwing didn't want to get over it. :)

OpenStudy (anonymous):

Thank you so much for helping!

OpenStudy (loser66):

;)

OpenStudy (anonymous):

I was in class at the time :DD